# CADDesigner: Conceptual Design of CAD Models Based on General-Purpose Agent

### v1

@@ Introduction
[…]
@@
Let the set of initial information input by the user be $\mathcal{I}(Text,Img)$, where $Text$ represents simple text description and $Img$ represents some sketchs. Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $F:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
@@
to check whether the model meets the user’s intent. When $Check(\mathcal{D},\mathcal{V})=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $Check(\mathcal{D},\mathcal{V})=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, ${}Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage. In the entire process, except for the user feedback links ($R$ and $Sat$), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of $State$ are independently completed by the CADDesigner).
@@
[…]
@@
[…]
@@
[…]
@@
where $\mathcal{G}\cap\mathcal{S}$ denotes the overlapping volume between the reference and generated models, and $\mathcal{G}\cup\mathcal{S}$ represents their combined volumetric union. A value of 111 indicates perfect alignment, while 0 signifies no overlap.
@@
[…]

### v2

@@ Introduction
This pipeline is augmented with a closed-loop correction mechanism: after generating visual feedback $\mathcal{V}$, the system invokes $T_{4}^{2}$ to obtain $(P,F)$. If $P=0$, the feedback $F$ is used to guide revised design generation via $T_{3}\circ T_{2}\circ T_{1}$. If $P=1$, the process terminates successfully. Thus, the loop continues until $T_{4}^{2}$ returns a termination signal.
@@
Let the set of initial information input by the user be $\mathcal{I}(Text,Img)$, where TextText represents simple text description and ImgImg represents some sketchs. Let $\mathcal{P}$ be the set of necessary parameters required for generating a CAD model. The process of the CADDesigner conducting detailed design can be expressed as $F:\mathcal{I}\to\mathcal{D}_{\text{detail}}$, where $\mathcal{D}_{\text{detail}}$ is detailed designs containing specific parameters, satisfying $\mathcal{P}\subseteq\mathcal{D}_{\text{detail}}$, which is accomplished by $T_{1}$. The user’s revision process for the detailed design is an iterative function:
@@
to check whether the model meets the user’s intent. When $Check(\mathcal{D},\mathcal{V})=0$, it returns visual feedback $F$ to the code generation stage to regenerate the code; when $Check(\mathcal{D},\mathcal{V})=1$, the result is delivered to the user for final judgment. The user’s satisfaction with the modeling result is determined by $Sat:\mathcal{M}\to\{0,1\}$. If $Sat(\mathcal{M})=1$, $\mathcal{M}$ is added to the case library $\mathcal{K}$, that is, $Case^{\prime}=Case\cup\{\mathcal{M}\}$; if $Sat(\mathcal{M})=0$, it re-enters the code generation stage. In the entire process, except for the user feedback links ($R$ and SatSat), the state transition function $Trans:State\to State^{\prime}$ (where the state transitions and judgments of StateState are independently completed by the CADDesigner).
@@
E.g., $s=\texttt{extrude\_rsolid(S, direction, distance)}$, in which extrude_rsolid is an Operation, S is an Object and $\texttt{direction},\texttt{distance}$ are Parameters.
@@
$\mathit{ErrMsg}=(\mathit{ErrCau},\ \mathit{ErrLoc},\ \mathit{CorrAct}),$
where $\mathit{ErrCau}$ denotes the root cause of failure, $\mathit{ErrLoc}$ indicates the precise location in the user script, and $\mathit{CorrAct}$ suggests one or more potential corrective actions.
@@
  • Function Annotations ($\mathit{Anno}$): semantic descriptions of commands, including parameter formats, return types, and usage notes;
  • Case Examples ($\mathit{Case}$): verified code snippets preserved as reference patterns.
@@
where $\mathcal{G}\cap\mathcal{S}$ denotes the overlapping volume between the reference and generated models, and $\mathcal{G}\cup\mathcal{S}$ represents their combined volumetric union. A value of 11 indicates perfect alignment, while 0 signifies no overlap.
@@
We sampled $1k$ instances in an even distribution of model complexity from the $8k$ test set of Text2CAD for comparison. For each instance, we used the abstract text description as input prompts. We then performed inference using their released weights and our method to generate results for metric calculation. As presented in Table 2 and Figure 3, the experimental results show that the model generated by our method achieves the optimal performance. Our method demonstrates the best prompt-result alignment, followed by CADCodeVerify and Text2CAD. Cadrille fails to generate valid outputs due to poor generalization, as it was trained on expert-level descriptions and cannot adapt to abstract-level inputs during inference.